# Quantum Models of Consciousness from a Quantum Information Science Perspective

### v2

@@ II Quantum Consciousness Emerging from the Microtubules Within Neurons
A key challenge for the Orch OR theory is whether quantum coherence can persist in the brain’s warm and noisy environment. Tegmark [46] estimated that decoherence in microtubules would occur in the order of $10^{-13}$ s, making quantum effects seemingly irrelevant. However, Hagan, Hameroff, and Tuszynski [47] recalculated decoherence times, accounting for dielectric properties, dipole interactions, and quantum shielding effects. Their revised estimates extended coherence times to $10^{-5}$$10^{-40}$ s, which aligns with neurophysiological processes. This time scale is crucial because it suggests that quantum coherence may be sustained long enough to influence neural processing and could play a role in the brain’s larger-scale functions.
@@
[…]
@@
The imaginary part $\Gamma$ of the complex eigenvalues ${\cal E}$ represents the decay width that determines the coupling of the extended excitonic state with the electromagnetic field. When the decay widths of all right eigenstates $|{\cal E}^{R}\rangle$ were plotted after being normalized by the single dipole decay width $\gamma$ in [61], it was realized that the superradiant state corresponds to the lowest excitonic state for the systems that have more than 12 spirals. It was also demonstrated that either full or a partial randomization of Trp dipole orientations destroys the superradiance, which in turn indicates that the superradiance depends on the particular order of these dipoles in microtubules.
@@
This type of probabilistic analysis of exciton delocalization within a microtubule segment could be further refined by incorporating established measures from quantum information theory. For instance, assume the microtubule system reaches thermal equilibrium at inverse temperature $\beta$, then, provided that the values of $\cal E$ are real or come in complex conjugate pairs with the same degeneracy [57, 58, 59, 60], the state of the system is given by
$\rho_{\text{th}}=\sum_{\cal E}\frac{e^{-\beta\cal E}}{Z}|\cal E^{R}\rangle\langle\cal E^{L}|,$ (3)
@@
[…]
@@ V Study of the Entanglement Preservation
In our approach, we treat each atom within a phosphate molecule as an identical spin having a magnitude of 1/2. In the context of a phosphate molecule, a cluster of five spins consists of a central spin (associated with the phosphorus atom) surrounded by buffer spins (corresponding to oxygen atoms). The buffer spins interact with individual thermal baths, while the central spin remains isolated from direct environmental interactions. Figure 1 from reference [29] illustrates this concept. When exposed to a magnetic field along the $z$-axis, an energy difference arises between the lower state (spin-up) and the upper state (spin-down), allowing for a two-level description. The buffer spins eventually reach thermal equilibrium with the environment at an inverse temperature $\beta$. In our study, we analyze the geometry of buffer networks characterized by different arrays of coupling constants, which are either $g$ or 00. These configurations can be represented using planar graphs. Specifically, for a given number of buffer spins, $N$, we explore all feasible buffer networks that can be embedded in a plane (Figure 2 from reference [29] illustrates the different planar graphs). We focus on two extreme cases: one where there is no connectivity among the buffer spins, and another where there is maximum connectivity within the buffer spin network.
@@
Table 1: Time, $t_{1}$, at which the logarithmic negativity between the two center spins is lower than $10^{-4}$. The left and right columns correspond to vanishing and maximal connectivities in the buffer network respectively at fixed $N$.
@@ VI.1 Hamiltonian Transformation
We now examine the system with a fully connected buffer network, and more specifically its Hamiltonian, from a different perspective, in an attempt to provide a theoretical explanation of better entanglement preservation in the fully connected geometry. By applying a unitary transformation to the Hamiltonian, we can transition from a fully connected graph of buffer spins to the other extreme, where the buffer spins do not interact with each other. These fictitious non-interacting buffer spins will be referred to as the dressed buffer spins and the new representation after the transformation as the dressed picture.
@@
$\Sigma_{+}$ $=(\sigma_{1}^{+},\dots,\sigma_{N}^{+}),$ (17)
$\Sigma_{-}$ $=(\sigma_{1}^{-},\dots,\sigma_{N}^{-}),$ (18)
@@
For $N=2,3,4$, this is equivalent to finding the eigenvalues and eigenvectors of the complete graph $A_{\text{com}}u_{j}=\mu_{j}u_{j}$. We readily obtain $\mu_{1}=\ldots=\mu_{N-1}=-1$ with multiplicity $N-1$, and $\mu_{N}=N-1$ with multiplicity 1. The corresponding eigenvectors are $u_{j}=\frac{1}{\sqrt{2}}\left(\ket{N}-\ket{j}\right)$ for $j=1,\ldots,N-1$, and $u_{N}=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\ket{j}.$ The matrix $D$ in (22) is diagonal and contains the frequencies of the dressed buffer spins, which, following our discussion on the eigenvalues of the complete graph and some trivial algebra for the geometry with $N=5$, are given by:
@@
The spin operator vectors are modified in the dressed picture as $\Sigma\rightarrow\tilde{\Sigma}$ with
$\tilde{\Sigma}=(\mathbf{1}\otimes U^{\dagger})\Sigma,$ (27)
which gives
$\tilde{\Sigma}=(\sigma_{1},\tilde{\sigma}_{2},\dots,\tilde{\sigma}_{N}).$ (28)
@@
As discussed before, $\sigma_{1}$, which represents the spin operator of the central spin, remains unchanged.
@@
Next, we rewrite the interaction between the central and the buffer spins in the dressed picture $\tilde{\Sigma}$, making use of the transformation,
$\sigma_{i}^{\pm}=\sum_{j=2}^{N+1}U_{ij}\tilde{\sigma_{j}}^{\pm}.$ (29)
@@
We can thus define a new vector of interactions $\tilde{v}=(\tilde{g}_{12},\dots,\tilde{g}_{1{N+1}})$. Note, however, that the coupling between the central spin and the buffer spins can be written as $G\sigma_{1}^{+}\sum_{j=2}^{N+1}\sigma_{j}^{-}+H.c.$, and in the case of a complete graph the sum $\sum_{j=2}^{N+1}\sigma_{j}^{-}$ corresponds to a single eigenvector of the graph ($u_{N}$ defined above, with eigenvalue $N-1$). In other words, for $N=2,3,4$, the central spin is coupled to a single dressed buffer spin with energy $\omega+(N-1)G$. Furthermore, the coupling constant is equal to $G\sqrt{N-1}$. Finally, for the case with $N=5$, the buffer spins do not form a perfectly complete graph, as the interaction between spin 2 and spin 6 is missing. However, a similar reasoning applies and it can be shown that the central spin is coupled to only two dressed buffer spins, as depicted in Figure 5 (column C). We remark that the decoupling of the central spin from most of the dressed buffer spins is due to the system Hamiltonian being highly symmetric.
@@
Let us finally address the coupling of the buffer spins with the local baths in the dressed picture. The local thermal dissipation channels are described by the dissipator (12). By applying the transformation (29), we obtain in the dressed picture the same dissipator with $\sigma_{i}$ replaced by the suitable linear combination of $\tilde{\sigma}_{j}$. In the dressed picture, the separate baths acting locally on each buffer spins become global baths [102, 103] acting collectively on the dressed buffer spins.
@@
In conclusion, through the change of variables (29) we have mapped the problem of a central spin coupled to a network of fully connected buffer spins to the problem of a central spin coupled to a collection of non-interacting fictitious buffer spins, which resembles the other model we consider in this work that displays worse entanglement preservation (see for instance Table 1). Contrary to the latter scenario, however, only one of the dressed buffer spins is coupled to the central spin, although through a stronger interaction and under the action of global baths. This may give us some hints of why the fully connected geometry better preserves entanglement, as we show in the following.
@@
Figure 5: The first column represents the number of buffer spins, $N$. Column A represents the non-interacting buffer spins. Column B shows the interacting buffer spins. Column C shows the dressed buffer spins after the Hamiltonian transformation which is represented by the curved arrows. In the original representation (column B), all spin couplings are identical. In the new representation, that is, the dressed picture (column C), the couplings vary as follows: $\tilde{g}_{12}=1.41G$ for $N=2$, $\tilde{g}_{12}=1.73G$ for $N=3$, $\tilde{g}_{12}=2G$ for $N=4$, and $\tilde{g}_{12}=-2.23G$, $\tilde{g}_{14}=0.21G$ for $N=5$.
@@ VI.2 Calculation for $N=2$
with $\tilde{g}_{12}=\sqrt{2}G$. It is worth noting that after diagonalization, the interactions between dressed buffer spins are suppressed.
@@ VI.3 Interpretation of the Results
In the non-interacting case, there are $N$ interactions between the central spin and buffer spins (column A in Figure 5). When interactions between buffer spins are introduced (column B in Figure 5), the interaction between the central spin and the dressed buffer spins is reduced to a single interaction for $N=2$ to $N=4$, and to two interactions for $N=5$ (as shown in column C in Figure 5). Using the dressed picture, we can more easily compare these scenarios and understand why interacting buffer spins offer better quantum information protection. In column B, where the buffer spins are maximally interacting, the interaction between the central spin and the buffer network (which is connected to the environment and responsible for the loss of quantum information) is significantly modified. While in the case of non-interacting buffer spins (column A) the central spin is weakly coupled to $N$ independent and locally dissipating buffer spins, if the buffer network is fully connected ($N=2,3,4$), then the central spin interacts with a single dressed buffer spin (column C), which is dissipating into different global baths in a more and more intense way as a function of $N$ (coupling constant going as $\sqrt{N}$).
@@
Focusing now on the dynamics of the central spin only, the difference between the two scenarios described above lies in a transition from a Markovian to a non-Markovian dynamics. Indeed, if dissipation comes from a collection of identical and weakly coupled systems, the dynamics have a more Markovian behavior corresponding to a roughly flat spectral density of the environment (this is an exact result if we consider $N\rightarrow\infty$ non-interacting buffer spins [101]). On the other hand, a stronger coupling with a single dressed buffer spin implies a more and more non-Markovian behavior, as the dressed spin filters the thermal white noise into Lorentzian noise. Furthermore, it is well-known that non-Markovian dynamics better preserve quantum information, as non-Markovianity induces a backflow of quantum information from the reservoir to the central spin [106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116]. It is thus reasonable to expect that the fully connected buffer network preserves entanglement during a longer time than the non-interacting network, as we indeed observe.
@@
For $N\geq 5$, it is physically impossible to construct a fully connected network, given the position of the central spin (see e.g., [29]), and for instance for $N=5$ the interaction between spin 2 and spin 6 is missing. So, the dressed buffer network does not reduce to a single interaction. Instead, the central spin is coupled to two non-interacting dressed buffer spins, while three more are left uncoupled. We can draw similar conclusions as for the case of perfectly fully connected buffer network, i.e., the dynamics will have a more non-Markovian behavior than in the case of non-interacting buffer network, leading to better entanglement preservation. This behavior, however, will be less pronounced than in the case of a fully connected network.
@@
Finally, we point out that in the tetrahedral geometry ($N=4$), there is a significant improvement in quantum information preservation (Figure 4). We observe that $N=4$ presents the strongest possible coupling between the central spin and a singled dressed buffer spin (Figure 5), as for $N\geq 5$ the fully connected geometry is not possible. Consequently, quantum information would be preserved longer due to the strongest possible non-Markovian behavior, arising in the case of $N=4$.
@@ VI.4 Numerical Verification
where $\|A\|_{1}$ denotes the trace norm of matrix $A$, which is the sum of the singular values of $A$. To verify non-Markovianity, we use the trace distance to compare two initially orthogonal states. The initial states of the central spins, |$\ket{\phi_{1}}$ and |$\ket{\phi_{2}}$, corresponding to $\rho_{1}$ and $\rho_{2}$ in the trace distance measure, are taken as
|$\ket{\phi_{1}}$ $=\frac{1}{\sqrt{2}}(\ket{0}+\ket{1});$ (36)
|$\ket{\phi_{2}}$ $=\frac{1}{\sqrt{2}}(\ket{0}-\ket{1}).$ (37)
@@
After evolving the two initial states, |$\ket{\phi_{1}}$ and |$\ket{\phi_{2}}$, we then compare them at each time instant using the trace norm distance. As shown in Figure 6, for $N=4$, we observe more pronounced oscillations, indicating higher non-Markovianity. This may explain our earlier findings and the enhanced protection observed in the tetrahedral geometry.

### v3

@@ II Quantum Consciousness Emerging from the Microtubules Within Neurons
A key challenge for the Orch OR theory is whether quantum coherence can persist in the brain’s warm and noisy environment. Tegmark [46] estimated that decoherence in microtubules would occur in the order of $10^{-13}$ s, making quantum effects seemingly irrelevant. However, Hagan, Hameroff, and Tuszynski [47] recalculated decoherence times, accounting for dielectric properties, dipole interactions, and quantum shielding effects. Their revised estimates extended coherence times to $10^{-5}$$10^{-4}$ s, which aligns with neurophysiological processes. This time scale is crucial because it suggests that quantum coherence may be sustained long enough to influence neural processing and could play a role in the brain’s larger-scale functions.
@@
$H_{\it eff}=H_{0}+\Delta-\frac{i}{2}G\,$, (1)
whose right and left eigenvectors do not form an orthogonal basis individually, but instead establish a biorthogonal basis collectively, a defining feature of non-Hermitian quantum systems [57, 58, 59, 60]. More precisely, given that $H_{\it eff}|{\cal E}^{R}\rangle={\cal E}|{\cal E}^{R}\rangle$ and $\langle{\cal E}^{L}|H_{\it eff}={\cal E}\langle{\cal E}^{L}|$, the eigenvectors satisfy the biorthogonality condition $\langle{\cal E^{\prime}}^{L}|{\cal E}^{R}\rangle=c_{\cal E}\,\delta_{\cal E^{\prime},\cal E}$, leading to the spectral decomposition $H_{\it eff}=\sum_{\cal E}{\cal E}|{\cal E}^{R}\rangle\langle{\cal E}^{L}|/c_{\cal E}$.
@@
[…]
@@
[…]
@@
In addition to enhancing the quantification of delocalization, the perspective of quantum information science could provide an opportunity to extend the non-Hermitian Hamiltonian approach to a quantum master equation approach. This would allow for modeling the complete dynamics of photoexcitation in microtubules, including the influence of the surrounding environment [53, 54, 55, 56]. The effective Hamiltonian $H_{\it eff}$ given in Equation (1), which characterizes the reduced photoexcitation dynamics, can be derived from a couple of extended master equations that include the electromagnetic field’s degrees of freedom [53, 54, 55, 56]. However, these master equations should also be extended to include thermal fluctuations of the surrounding environment.
@@ V Study of the Entanglement Preservation
In our approach, we treat each atom within a phosphate molecule as an identical spin having a magnitude of 1/2. In the context of a phosphate molecule, a cluster of five spins consists of a central spin (associated with the phosphorus atom) surrounded by buffer spins (corresponding to oxygen atoms). The buffer spins interact with individual thermal baths, while the central spin remains isolated from direct environmental interactions. Figure 1 from reference [29] illustrates this concept. When exposed to a magnetic field along the $z$-axis, an energy difference arises between the lower state (spin-up) and the upper state (spin-down), allowing for a two-level description. The buffer spins eventually reach thermal equilibrium with the environment at an inverse temperature $\beta$. In our study, we analyze the geometry of buffer networks characterized by different arrays of coupling constants, which are either $g$ or 0. These configurations can be represented using planar graphs. Specifically, for a given number of buffer spins, $N$, we explore all feasible buffer networks that can be embedded in a plane (Figure 2 from reference [29] illustrates the different planar graphs). We focus on two extreme cases: one where there is no connectivity among the buffer spins, and another where there is maximum connectivity within the buffer spin network.
@@
Table 1: Time, $t_{1}$, at which the logarithmic negativity between the two center spins is lower than $10^{-4}$. The left and right columns correspond to vanishing and maximal connectivities in the buffer network respectively at fixed NN.
@@ VI.1 Hamiltonian Transformation
We now examine the system with a fully connected buffer network, and more specifically its Hamiltonian, from a different perspective, in an attempt to provide a theoretical explanation of better entanglement preservation in the fully connected geometry. By applying a unitary transformation to the Hamiltonian, we can transition from a fully connected graph of buffer spins to the other extreme, where the buffer spins are mapped to a set of non-interacting collective modes. These fictitious non-interacting modes will be referred to as the dressed buffer modes and the new representation after the transformation as the dressed picture.
@@
$\Sigma_{+}$ $=(\sigma_{1}^{+},\dots,\sigma_{N+1}^{+}),$ (17)
$\Sigma_{-}$ $=(\sigma_{1}^{-},\dots,\sigma_{N+1}^{-}),$ (18)
@@
[…]
@@
The spin operators of the buffer spins are mapped to some new operators associated with a set of dressed collective excitations. It is worth remarking that, in general, the dressed modes are not spins anymore, and the dressed operators in general do not satisfy the spin commutation rules. With this in mind, we refer to $\tilde{\sigma}_{j}^{-}$ and $\tilde{\sigma}_{j}^{+}$ with $j=2,\ldots,N+1$ as respectively the lowering and raising operators of the dressed buffer modes. The relation between the set of operators before and after the diagonalization of $H^{\prime}$ is
$\tilde{\Sigma}_{+}^{T}=(\mathbf{1}\otimes U^{T})\Sigma_{+}^{T},\qquad\tilde{\Sigma}_{-}^{T}=(\mathbf{1}\otimes U^{\dagger})\Sigma_{-}^{T},$ (27)
with the identity acting on the first central spin, and
$\tilde{\Sigma}_{\pm}=(\sigma_{1}^{\pm},\tilde{\sigma}_{2}^{\pm},\dots,\tilde{\sigma}_{N}^{\pm}).$ (28)
@@
As discussed before, $\sigma_{1}^{\pm}$, which represent the spin operators of the central spin, remain unchanged.
@@
Next, we rewrite the interaction between the central and the buffer spins in the dressed picture $\tilde{\Sigma}$, making use of the transformation
$\sigma_{i}^{+}=\sum_{j=2}^{N+1}U_{ij}^{*}\tilde{\sigma}_{j}^{+},\quad\sigma_{i}^{-}=\sum_{j=2}^{N+1}U_{ij}\tilde{\sigma}_{j}^{-},\quad\text{for }i=2,\ldots,N+1.$ (29)
We can thus define a new vector of interactions $\tilde{v}=(\tilde{g}_{12},\dots,\tilde{g}_{1{N+1}})$. Note, however, that the coupling between the central spin and the buffer spins can be written as $G\sigma_{1}^{+}\sum_{j=2}^{N+1}\sigma_{j}^{-}+H.c.$, and in the case of a complete graph the sum $\sum_{j=2}^{N+1}\sigma_{j}^{-}$ corresponds to a single eigenvector of the graph ($u_{N}$ defined above, with eigenvalue N1N-1). Moreover, this may also be thought of as the operator of a collective spin acting on all buffer spins and decoupled from all other buffer modes. The lowering and raising operators of the dressed buffer mode associated with the collective spin are properly renormalized by $1/\sqrt{N-1}$, so that their coupling to the central spin grows as $\sqrt{N}$. In addition, the frequency of this dressed mode grows as $N$.
@@
Finally, for the case with $N=5$, the buffer spins do not form a perfectly complete graph, as the interaction between spin 2 and spin 6 is missing. However, a similar reasoning applies and it can be shown that the central spin is coupled to only two dressed buffer modes, as depicted in Figure 5 (column C). In this case, however, these collective modes cannot be associated with collective spins. We remark that the decoupling of the central spin from most of the dressed buffer modes is due to the system Hamiltonian being highly symmetric.
@@
Let us finally address the coupling of the buffer spins with the local baths in the dressed picture. The local thermal dissipation channels are described by the dissipator (12). By applying the transformation (29), we obtain in the dressed picture the same dissipator with $\sigma_{i}$ replaced by the suitable linear combination of $\tilde{\sigma}_{j}$. In the dressed picture, the separate baths acting locally on each buffer spins become global baths [102, 103] acting collectively on the dressed buffer modes.
@@
In conclusion, through the change of variables (29) we have mapped the problem of a central spin coupled to a network of fully connected buffer spins to the problem of a central spin coupled to a collection of non-interacting buffer modes. For $N=2,3,4$, the central spin interacts with a single dressed buffer mode that may be associated with a collective spin, while all other dressed modes are decoupled from the central spin. This scenario resembles the other model we consider in this work that displays worse entanglement preservation (see for instance Table 1). Contrary to the latter scenario, however, only one collective “spin-like” buffer mode is coupled to the central spin, although through a stronger interaction and under the action of global baths. This may give us some hints of why the fully connected geometry better preserves entanglement, as we show in the following. For $N=5$ this argument is weaker, as the central spin interacts with two generic collective dressed modes.
@@
Figure 5: The first column represents the number of buffer spins, NN. Column A represents the non-interacting buffer spins. Column B shows the interacting buffer spins. Column C shows the dressed buffer modes after the Hamiltonian transformation which is represented by the curved arrows. For $N=2,3,4$, the dressed mode $\tilde{2}$ can be associated with a collective spin. In the original representation (column B), all spin couplings are identical. In the new representation, that is, the dressed picture (column C), the couplings vary as follows: $\tilde{g}_{12}=1.41G$ for $N=2$, $\tilde{g}_{12}=1.73G$ for $N=3$, $\tilde{g}_{12}=2G$ for $N=4$, and $\tilde{g}_{12}=-2.23G$, $\tilde{g}_{14}=0.21G$ for $N=5$.
@@ VI.2 Calculation for $N=2$
with $\tilde{g}_{12}=\sqrt{2}G$. It is worth noting that after diagonalization, the interactions between dressed buffer modes are suppressed.
@@ VI.3 Interpretation of the Results
In the non-interacting case, there are $N$ interactions between the central spin and buffer spins (column A in Figure 5). When interactions between buffer spins are introduced (column B in Figure 5), the interaction between the central spin and the dressed buffer modes is reduced to a single interaction for $N=2$ to $N=4$, and to two interactions for $N=5$ (as shown in column C in Figure 5). Using the dressed picture, we can more easily compare these scenarios and understand why interacting buffer spins offer better quantum information protection. In column B, where the buffer spins are maximally interacting, the interaction between the central spin and the buffer network (which is connected to the environment and responsible for the loss of quantum information) is significantly modified. While in the case of non-interacting buffer spins (column A) the central spin is weakly coupled to $N$ independent and locally dissipating buffer spins, if the buffer network is fully connected ($N=2,3,4$), then the central spin interacts with a single dressed collective buffer modes (column C), which may be associated with a collective spin, in a more and more intense way as a function of $N$ (coupling constant going as $\sqrt{N}$). This mode is in turn dissipating into different global baths.
@@
Focusing now on the dynamics of the central spin only, the difference between the two scenarios described above lies in a transition from a Markovian to a non-Markovian dynamics. Indeed, if dissipation comes from a collection of identical and weakly coupled systems, the dynamics have a more Markovian behavior corresponding to a roughly flat spectral density of the environment (this is an exact result if we consider $N\rightarrow\infty$ non-interacting buffer spins [101]). On the other hand, a stronger coupling with a single dressed buffer mode implies a more and more non-Markovian behavior, as the dressed mode filters the thermal white noise into Lorentzian noise. Furthermore, it is well-known that non-Markovian dynamics better preserve quantum information, as non-Markovianity induces a backflow of quantum information from the reservoir to the central spin [106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116]. It is thus reasonable to expect that the fully connected buffer network preserves entanglement during a longer time than the non-interacting network, as we indeed observe.
@@
For $N\geq 5$, it is physically impossible to construct a fully connected network, given the position of the central spin (see e.g., [29]), and for instance for $N=5$ the interaction between spin 2 and spin 6 is missing. So, the dressed buffer network does not reduce to a single interaction. Instead, the central spin is coupled to two non-interacting dressed buffer modes, while three more are left uncoupled. Although the analogy with the case of uncoupled buffer spins is partially lost, we might still draw similar conclusions as for the case of perfectly fully connected buffer network, i.e., the dynamics will have a more non-Markovian behavior than in the case of non-interacting buffer network, leading to better entanglement preservation. This behavior, however, will be less pronounced than in the case of a fully connected network.
@@
Finally, we point out that in the tetrahedral geometry ($N=4$), there is a significant improvement in quantum information preservation (Figure 4). We observe that $N=4$ presents the strongest possible coupling between the central spin and a singled dressed buffer mode (Figure 5), as for $N\geq 5$ the fully connected geometry is not possible. Consequently, quantum information would be preserved longer due to the strongest possible non-Markovian behavior, arising in the case of $N=4$.
@@ VI.4 Numerical Verification
[…]
@@
[…]